# Optimal Transport for $\epsilon$-Contaminated Credal Sets

### v1

@@ Abstract
We provide a version for lower probabilities of Monge’s and Kantorovich’s optimal transport problems. We show that, when the lower probabilities are the lower envelopes of $\epsilon$-contaminated sets, then our version of Monge’s, and a restricted version of our Kantorovich’s problems, coincide with their respective classical versions. We also give sufficient conditions for the existence of our version of Kantorovich’s optimal plan, and for the two problems to be equivalent. As a byproduct, we show that for $\epsilon$-contaminations the lower probability versions of Monge’s and Kantorovich’s optimal transport problems need not coincide. The applications of our results to Machine Learning and Artificial Intelligence are also discussed.
@@ 1 Introduction
The concept of stochasticity is pervasive in modern-day artificial intelligence (AI) and machine learning (ML), allowing to capture the lack of determinism that underpins virtually all interesting applications, ranging from the medical domain Stutz et al. (2023) to trajectory prediction of ballistic missiles (Ji et al., 2022).
@@
Via a classical measure-theoretic argument (Rudin, 1987), we can obtain $P_{2}$ as the pushforward measure of $P_{1}$ via $T$,
$P_{2}(\xi_{2}\in B)$ $=P_{1}(T(\xi_{1})\in B)=P_{1}(\xi_{1}\in T^{-1}(B)),$
where $B$ is an arbitrary subset of the space that $\xi_{2}$ take values on. Then, we can write $P_{2}\equiv T_{\#}P_{1}\coloneqq P_{1}\circ T^{-1}$, so that $P_{2}$ is indeed the pushforward measure $T_{\#}P_{1}$ of $P_{1}$ via $T$.
An interesting question we may ask ourselves at this point is whether we can turn the problem around. Given $P_{1}$ and $P_{2}$, there are many functions $T$ that push $P_{1}$ to $P_{2}$. Is there an “optimal” one, that is, one that makes the transformation from $P_{1}$ to $P_{2}$ as efficient (i.e. less expensive) as possible? This goal is similar to one that Napoleonic engineers were assigned by Napoleon himself: they were tasked to find the cheapest way of transporting iron ore from the mines to the factories (Villani, 2003).
To find such an optimal $T$, in the late 1700s Gaspard Monge suggested the following optimization problem
$\inf\left\{{\int c(\xi_{1},T(\xi_{1}))P_{1}(\text{d}\xi_{1}):T_{\#}P_{1}=P_{2}}\right\},$
where $c$ is a cost function gauging the cost of moving one unit of probability mass from $\xi_{1}$ to $T(\xi_{1})$, which in turn gives us a measure of the efficiency of “moving probability bits” from $P_{1}$ to $P_{2}=T_{\#}P_{1}$.
Alas, an optimal solution $T$ to this problem may not exist, for example when $P_{1}$ is a Dirac measure, but $P_{2}$ is not. Fortunately, though, Leonid Kantorovich came up with an equivalent formulation of the problem that, under mild conditions, is guaranteed to be well-posed. His expression is the following
$\inf\left\{{\int c(\xi_{1},\xi_{2})\alpha(\text{d}(\xi_{1},\xi_{2})):\alpha\in\Gamma(P_{1},P_{2})}\right\},$
where $\Gamma(P_{1},P_{2})$ is the set of all joint probability measures whose marginals are $P_{1}$ and $P_{2}$. Instead of looking for the most efficient transportation map $T$ from $\xi_{1}$ to $\xi_{2}$, it seeks the “cheapest” transportation plan $\alpha$ between the distributions $P_{1}$ and $P_{2}$. Kantorovich’s problem is solvable via linear programming, and it coincides with Monge’s problem when $\alpha(\text{d}(\xi_{1},\xi_{2}))=P_{1}(\text{d}\xi_{1})\delta_{\xi_{2}=T(\xi_{1})}$, where the latter is the Dirac measure for the event $\{\xi_{2}=T(\xi_{1})\}$ (Thorpe, 2018, Section 1.2).
@@
$W_{p}(P_{1},P_{2})\coloneqq\left[\inf_{\alpha\in\Gamma(P_{1},P_{2})}\int d(\xi_{1},\xi_{2})^{p}\alpha(\text{d}(\xi_{1},\xi_{2}))\right]^{\frac{1}{p}}$
@@
Contributions. In this paper, we ask ourselves what do Monge’s and Kantorovich’s problems look like, when instead of transporting probability measures, we transport lower probabilities, their imprecise counterpart that allow to describe the ambiguity faced by the scholar around the true data generating processes (Walley, 1991; Augustin et al., 2014; Troffaes and de Cooman, 2014). We give the first ever (to the best of our knowledge) definitions of Monge’s and Kantorovich’s problems for lower probabilities, and then we focus our attention on sets of probabilities $\mathcal{M}(\underline{P})$ (called credal sets) that are completely characterized by their lower envelope $\underline{P}$ (that is a lower probability). This means that the whole set $\mathcal{M}(\underline{P})$ can be reconstructed by simply looking at lower probability $\underline{P}=\inf_{P\in\mathcal{M}(\underline{P})}P$. We show that, for a particular class of such sets of probabilities, called the $\epsilon$-contamination class (which we define in (4)), the optimal Monge’s map (when it exists) and (a version of) the optimal Kantorovich’s transportation plan coincide with their classical counterparts. This is an important result, as it promises to be fraught with fruitful consequences for many possible applications. We go on giving sufficient conditions (i) for the existence of the lower probability version of Kantorovich’s optimal plan, and (ii) for the two problems formulations to coincide. A byproduct of the latter is that, in general, the lower probability versions of Monge’s and Kantorovich’s problems need not coincide, a fact that we conjecture holds not only for $\epsilon$-contaminations, but for all credal sets.
@@
Motivation and Related Work. Besides being interested in these results for their own mathematical beauty, our motivations to study them stem from the field of Imprecise Probabilistic Machine Learning (IPML) (Caprio et al., 2023; Caprio and Gong, 2023; Caprio and Mukherjee, 2023; Hüllermeier and Waegeman, 2021; Caprio et al., 2024; Denœux and Zouhal, 2001; Destercke et al., 2008; Lu et al., 2024; Zaffalon, 2002). Credal Machine Learning (CML), a subfield of IPML, devotes itself to developing ML theory and methods working with credal sets. Our findings in this paper can be immediately applied to CML in at least three different contexts.
@@
First, they can be used to define new uncertainty measures enjoying the axiomatic desiderata in Abellán and Klir (2005); Jiroušek and Shenoy (2018); Sale et al. (2023) based on Hausdorff-type distances between sets of probabilities, which would extend the results in Sale et al. (2024) to credal sets.
@@
A second immediate application is robust Hypotheses Testing (HT) (Acharya et al., 2015; Gao et al., 2018; Mortier et al., 2023; Liu and Briol, 2024; Chau et al., 2024), where a test statistic based on the optimal transport cost between lower probabilities characterizing credal sets could be used to test whether the true data generating process – that produces the data accruing to the phenomenon of interest – belongs to either of the credal sets. In HT notation, $H_{0}:P^{\text{true}}$ belongs to $\mathcal{M}(\underline{P}_{1})$, versus $H_{1}:P^{\text{true}}$ belongs to $\mathcal{M}(\underline{P}_{2})$.
@@
Finally, our findings can be seminal in starting a new field of inquiry that analyzes optimal transport problems involving credal sets in the setting of credal ergodic theory, extending e.g. Lopes and Mengue (2012); Kabir and Lee (2020); Chen and Li (2020) to the credal framework. This would greatly benefit computer vision (and especially the theory of convolutional autoencoders (Yu et al., 2023)) via the imprecise Markov semigroup approach of Caprio (2024).
@@
Structure of the Paper. The paper is arranged as follows. Section 2 gives the necessary background on credal sets. Our results pertaining the lower probability version of Monge’s and Kantorovich’s problems are presented in Sections 3 and 3.1, respectively. Section 4 concludes our work.
@@ 2 Background
In this work, we focus on a particular type of credal sets (convex and weak-closed sets of probabilities (Levi, 1980)), [^1] We recall the definition of weak topology in the proof of Lemma 2. that is, those that economists and operations researchers call cores (of an exact capacity) (Cerreia-Vioglio et al., 2015; Caprio and Mukherjee, 2023; Miranda and Montes, 2023).
Given a capacity of interest – in this paper, it will always be a lower probability $\underline{P}$, i.e. a set function on the $\sigma$-algebra on interest, mapping in $[0,1]$, which is the lower envelope of a weak-compact set (Cerreia-Vioglio et al., 2015, Section 2.1.(viii)) – on a generic measurable space $(\mathcal{X},\mathcal{F})$, the core is defined as
$\mathcal{M}(\underline{P})\coloneqq\{P\in\Delta_{\mathcal{X}}:P(A)\geq\underline{P}(A)\text{, }\forall A\in\mathcal{F}\},$ (1)
where $\Delta_{\mathcal{X}}$ denotes the set of countably additive probabilities on $(\mathcal{X},\mathcal{F})$.
We focus on cores for two main reasons. First, in general we have that the convex hull of a finite set of countably additive probabilities on $\mathcal{X}$ is a proper subset of the core of the lower probability associated with that set, see e.g. (Amarante and Maccheroni, 2006, Example 1) and (Amarante et al., 2006, Examples 6,7,8). That is, given $\{P_{k}\}_{k=1}^{K}\subset\Delta_{\mathcal{X}}$, $K<\infty$, we have that $\text{CH}(\{P_{k}\}_{k=1}^{K})\subset\text{core}(\underline{P})$, where CH denotes the convex hull operator, and $\underline{P}(A)=\inf_{P\in\text{CH}(\{P_{k}\}_{k=1}^{K})}P(A)$, for all $A\in\mathcal{F}$. Hence, focusing on the core gives us more generality.
Second, the core is uniquely identified by its lower probability (Gong and Meng, 2021). To see this, notice that by knowing $\underline{P}$, we can reconstruct the set by simply considering all countably additive probability measures on $\mathcal{X}$ that set-wise dominate $\underline{P}$.
Before giving the main results of this paper, we need to introduce the concepts of pushforward lower probability (PLP) and of Choquet integral.
@@
Given two measurable spaces $(\mathcal{X},\mathcal{F})$ and $(\mathcal{Y},\mathcal{G})$, a measurable mapping $T:\mathcal{X}\rightarrow\mathcal{Y}$, and a lower probability $\underline{P}=\inf_{P\in\mathcal{M}(\underline{P})}P$, the pushforward of $\underline{P}$ is the (set) function $T_{\#}\underline{P}:\mathcal{G}\rightarrow[0,1]$ such that
@@
Proof.
[…]
$\text{core}(T_{\#}\underline{P})\coloneqq\{\mu\in\text{ba}(\mathcal{G}):\mu(B)\geq T_{\#}\underline{P}(B)\text{, }\forall B\in\mathcal{G}\text{, and }\mu(\mathcal{Y})=T_{\#}\underline{P}(\mathcal{Y})\},$
where $\text{ba}(\mathcal{G})$ is the vector spaces of all bounded charges (signed, finitely additive measures) on $\mathcal{G}$.
Notice that $\text{core}(T_{\#}\underline{P})$ is nonempty because it contains the pushforward measures of the elements of $\mathcal{M}(\underline{P})$ in (1) through map $T$. In formulas, $P\in\mathcal{M}(\underline{P})\implies T_{\#}P\in\text{core}(T_{\#}\underline{P})$. To see that this is the case, recall that countable additivity implies finite additivity, and that $\mathcal{M}(\underline{P})$ contains all the countably additive probabilities on $\mathcal{X}$ that set-wise dominate $\underline{P}$. More formally, pick any $\tilde{B}\in\mathcal{G}$ and any $P\in\mathcal{M}(\underline{P})$. Let $\tilde{A}=T^{-1}(\tilde{B})$. Then, by (1),
$T_{\#}P(\tilde{B})=P(T^{-1}(\tilde{B}))=P(\tilde{A})\geq\underline{P}(\tilde{A})=\underline{P}(T^{-1}(\tilde{B}))=T_{\#}\underline{P}(\tilde{B}).$
But then $T_{\#}P\in\text{core}(T_{\#}\underline{P})$, which shows that $\text{core}(T_{\#}\underline{P})\neq\emptyset$. In turn, by Marinacci and Montrucchio (2004, Proposition 3), we have that $\text{core}(T_{\#}\underline{P})$ is weak-compact.
Now, notice that $\mathcal{M}(T_{\#}\underline{P})\coloneqq\{Q\in\Delta_{\mathcal{Y}}:Q(B)\geq T_{\#}\underline{P}(B)\text{, }\forall B\in\mathcal{G}\}$ is a proper subset of $\text{core}(T_{\#}\underline{P})$, since $\mathcal{M}(T_{\#}\underline{P})$ only considers the countably additive probabilities that set-wise dominate $T_{\#}\underline{P}$.
Let now $(Q_{\alpha})_{\alpha\in\mathcal{I}}$ be a net in $\mathcal{M}(T_{\#}\underline{P})$ that weak-converges to $Q\in\Delta_{\mathcal{Y}}$. Here $\mathcal{I}$ is a generic index set. This means that $Q_{\alpha}(B)\rightarrow Q(B)$, for all $B\in\mathcal{G}$. That is, pick any $B\in\mathcal{G}$; then,
Equation (2) implies that, for all $Q_{\alpha}\succeq Q_{\tilde{\alpha}_{\epsilon}}$, we have that $Q(B)>Q_{\alpha}(B)-\epsilon\geq T_{\#}\underline{P}(B)-\epsilon$. Letting $\epsilon\rightarrow 0$, this implies that $Q(B)\geq T_{\#}\underline{P}(B)$. But $B$ was chosen arbitrarily in $\mathcal{G}$, and so $Q\in\mathcal{M}(T_{\#}\underline{P})$. Hence, $\mathcal{M}(T_{\#}\underline{P})$ is weak sequentially closed, and therefore it is weak closed. Being a weak closed subset of a weak compact space, we can conclude that $\mathcal{M}(T_{\#}\underline{P})$ is weak compact itself.
By Cerreia-Vioglio et al. (2015, Section 2.1.(viii)), we know that a (set) function $\nu:\mathcal{G}\rightarrow[0,1]$ is a lower probability if and only if there exists a weak-compact set $\mathcal{M}\subseteq\Delta_{\mathcal{Y}}$ such that $\nu(B)=\min_{Q\in\mathcal{M}}Q(B)$, for all $B\in\mathcal{G}$. Letting $\mathcal{M}\equiv\mathcal{M}(T_{\#}\underline{P})$ and $\nu\equiv T_{\#}\underline{P}$ concludes the proof. ∎
Lemma 2 entails that $T_{\#}\underline{P}$ is a superadditive version of a pushforward probability measure (Walley, 1991, Section 1.6.4). We now introduce Choquet integrals (Choquet, 1954),(Troffaes and de Cooman, 2014, Section C.2).
@@
$f\mapsto\int_{\mathcal{Z}}f\text{d}\underline{P}\coloneqq\int_{0}^{\infty}\underline{P}(\{f^{+}\geq t\})\text{d}t-\int_{0}^{\infty}\left[\underline{P}(\mathcal{Z})-\underline{P}(\{f^{-}\geq t\}^{c})\right]\text{d}t$ (3)
called the Choquet integral of $f$ with respect to $\underline{P}$, provided that the difference on the right-hand side is well defined. There, $f^{+}\coloneqq 0\vee f$, $f^{-}\coloneqq-(0\wedge f)$, and the integrals are (improper) Riemannian integrals. [^3] For a primer on Riemannian integrals, see Troffaes and de Cooman (2014, Section C.1).
@@
Using the same notation as Definition 3, suppose the Choquet integral $\int_{\mathcal{Z}}f\text{d}\underline{P}$ of $f$ with respect to $\underline{P}$ exists. Then,
$\int_{\mathcal{Z}}f\text{d}\underline{P}=\int_{0}^{\infty}\underline{P}(\{f\geq t\})\text{d}t+\int_{-\infty}^{0}\left[\underline{P}(\{f\geq t\})-\underline{P}(\mathcal{Z})\right]\text{d}t.$
In addition, if $f$ is bounded, then it is Choquet integrable with respect to $\underline{P}$, that is, its Choquet integral $\int_{\mathcal{Z}}f\text{d}\underline{P}$ exists.

### v3

@@ Abstract
We present generalized versions of Monge’s and Kantorovich’s optimal transport problems with the probabilities being transported replaced by lower probabilities. We show that, when the lower probabilities are the lower envelopes of $\epsilon$-contaminated sets, then our version of Monge’s, and a restricted version of our Kantorovich’s problems, coincide with their respective classical versions. We also give sufficient conditions for the existence of our version of Kantorovich’s optimal plan, and for the two problems to be equivalent. As a byproduct, we show that for $\epsilon$-contaminations the lower probability versions of Monge’s and Kantorovich’s optimal transport problems need not coincide. The applications of our results to Machine Learning and Artificial Intelligence are also discussed.
@@ 1 Introduction
The concept of stochasticity is pervasive in modern-day artificial intelligence (AI) and machine learning (ML), allowing to capture the lack of determinism that underpins virtually all interesting applications, ranging from the medical domain (Stutz et al., 2023; Caprio et al., 2025) to trajectory prediction of ballistic missiles (Ji et al., 2022).
@@
From a classical measure-theoretic argument (Rudin, 1987), we can obtain $P_{2}$ as the pushforward measure of $P_{1}$ via $T$, $P_{2}(\xi_{2}\in B)=P_{1}(T(\xi_{1})\in B)=P_{1}(\xi_{1}\in T^{-1}(B))$, where $B$ is an arbitrary subset of the space that $\xi_{2}$ take values on. Then, we can write $P_{2}\equiv T_{\#}P_{1}\coloneqq P_{1}\circ T^{-1}$, so that $P_{2}$ is indeed the pushforward measure $T_{\#}P_{1}$ of $P_{1}$ via $T$.
An interesting question we may ask ourselves at this point is whether we can turn the problem around. Given $P_{1}$ and $P_{2}$, there are many functions $T$ that push $P_{1}$ to $P_{2}$. Is there an “optimal” one, that is, one that makes the transformation from $P_{1}$ to $P_{2}$ as efficient (i.e. less expensive) as possible? This question, which lays at the heart of the field of Optimal Transport (OT), is similar to one that Napoleonic engineers were asked by Napoleon himself. They were tasked to find the cheapest way of transporting iron ore from the mines to the factories (Villani, 2003).
To find such an optimal $T$, in the late 1700s Gaspard Monge suggested the following optimization problem,
$\operatorname*{arg\,inf}\left\{{\int c(\xi_{1},T(\xi_{1}))P_{1}(\text{d}\xi_{1}):T_{\#}P_{1}=P_{2}}\right\}.$
It seeks to find the function $T$ that makes transporting the probability mass encoded in $P_{1}$ to that encoded in its pushforward via $T$, $P_{2}=T_{\#}P_{1}$, as “cheap” as possible. The latter is gauged by considering a cost function $c$ that gives us the cost of moving one unit of probability mass from $\xi_{1}$ to $\xi_{2}=T(\xi_{1})$. In other words, $c$ gives us a measure of the efficiency of “moving probability bits” from $P_{1}$ to $P_{2}=T_{\#}P_{1}$.
Alas, an optimal solution $T$ to this problem may not exist (Thorpe, 2018, Section 1.2). Fortunately, though, Leonid Kantorovich came up with an equivalent formulation of the problem that, under mild conditions, is guaranteed to be well-posed. His expression is the following
$\operatorname*{arg\,inf}\left\{{\int c(\xi_{1},\xi_{2})\text{d}\alpha(\xi_{1},\xi_{2}):\alpha\in\Gamma(P_{1},P_{2})}\right\},$
where $\Gamma(P_{1},P_{2})$ is the set of all joint probability measures whose marginals are $P_{1}$ and $P_{2}$. Instead of looking for the most efficient transportation map $T$ from $\xi_{1}$ to $\xi_{2}$, it seeks the “cheapest” transportation plan $\alpha$ between the distributions $P_{1}$ and $P_{2}$. The relationship between the optimal transportation plan $\alpha$ and the theory of copulas (Nelsen, 2006) [^1] Recall that a copula is a multivariate cumulative distribution function, for which the marginal probability distribution of each variable is Uniform on the interval $[0,1]$. was studied e.g. in Chi et al. (2022); Liu et al. (2023).
@@
$W_{p}(P_{1},P_{2})\coloneqq\left[\inf_{\alpha\in\Gamma(P_{1},P_{2})}\int d(\xi_{1},\xi_{2})^{p}\text{d}\alpha(\xi_{1},\xi_{2})\right]^{\frac{1}{p}}$
@@
Contributions. In this paper, we ask ourselves:
Question 1: What do Monge’s and Kantorovich’s problems look like, when instead of transporting probability measures, we transport lower probabilities?
@@
Lower probabilities are the imprecise counterpart of classical probabilities that allow to describe the ambiguity faced by the scholar around the true data generating processes (Walley, 1991; Augustin et al., 2014; Troffaes and de Cooman, 2014). We give the first (to the best of our knowledge) definitions of Monge’s and Kantorovich’s problems for lower probabilities, and then we focus our attention on sets of probabilities $\mathcal{M}(\underline{P})$ (called cores of a coherent lower probability, a special type of credal sets) that are completely characterized by their lower envelope $\underline{P}$ (that is a lower probability). [^2] In general, credal sets are in one-to-one correspondence with lower prevision functionals (Troffaes and de Cooman, 2014). This means that the whole set $\mathcal{M}(\underline{P})$ can be reconstructed by simply looking at lower probability $\underline{P}=\inf_{P\in\mathcal{M}(\underline{P})}P$. A pictorial representation of our endeavor is given in Figure 1.
It is necessary to mention that we are not the first to extend the study of OT beyond classical probability theory. Lorenzini et al. (2024); Nguyen (2023) do so for belief functions and random sets, and Rachev and Olkin (1999); Gal and Niculescu (2019); Torra (2023) do so for capacities and non-additive measures.
Question 2: Is there a class of credal sets completely characterized by their lower probabilities (LPs), for which the LP versions of Monge’s and Kantorovich’s problems coincide with their classical counterparts?
We show that, for the class of $\epsilon$-contaminated credal sets (which we introduce in (3)), the answer to Question 2 is positive. [^3] What we mean is that solving the LP version of Monge’s and Kantorovich’s problems is equivalent to solving their classical versions. The solutions, then, will trivially coincide. This is an important result, as it promises to be fraught with fruitful consequences for many possible applications. We also give sufficient conditions (i) for the existence of the lower probability version of Kantorovich’s optimal plan, and (ii) for the two problems formulations to coincide. A byproduct of the latter is that, in general, the lower probability versions of Monge’s and Kantorovich’s problems need not coincide.
Figure 1: Top: the optimal transport map $T$ between two bell-shaped distributions $P$ and $Q$ on $\mathbb{R}$. Bottom: the optimal transport map $T$ between lower probabilities $\underline{P}$ and $\underline{Q}$ (both depicted as black brushstrokes) that completely characterize credal sets $\mathcal{P}$ and $\mathcal{Q}$ of probabilities on $\mathbb{R}$. The colored distributions are elements of the respective credal sets.
@@
Motivation and Related Work. Besides being interested in these results for their own mathematical beauty, our motivations to study them stem from the field of Imprecise Probabilistic Machine Learning (IPML) (Caprio et al., 2024a; Caprio and Gong, 2023; Caprio and Mukherjee, 2023; Hüllermeier and Waegeman, 2021; Caprio et al., 2024b; Denœux and Zouhal, 2001; Destercke et al., 2008; Lu et al., 2024; Zaffalon, 2002). Credal Machine Learning (CML), a subfield of IPML, devotes itself to developing ML theory and methods working with credal sets. Our findings in this paper can be immediately applied to CML in at least three different contexts.
@@
First, they can be used to define new uncertainty measures enjoying the axiomatic desiderata in Abellán and Klir (2005); Jiroušek and Shenoy (2018); Sale et al. (2023) based on Hausdorff-type distances between sets of probabilities, which would extend the results in Sale et al. (2024) to credal sets. To be more specific, the Hausdorff distance based on the Wasserstein metric between two credal sets could be expressed as the Wasserstein distance between their lower probabilities. The latter is intimately related to the Kantorovich’s OT problem that we define in this paper.
@@
A second immediate application is robust Hypotheses Testing (HT) (Augustin, 2002; Acharya et al., 2015; Gao et al., 2018; Mortier et al., 2023; Liu and Briol, 2024; Chau et al., 2025; Huber and Strassen, 1973), where a test statistic based on the optimal transport cost between lower probabilities characterizing credal sets could be used to test whether the true data generating process – that produces the data accruing to the phenomenon of interest – belongs to either of the credal sets. In HT notation, $H_{0}:P^{\text{true}}$ belongs to $\mathcal{M}(\underline{P}_{1})$, versus $H_{1}:P^{\text{true}}$ belongs to $\mathcal{M}(\underline{P}_{2})$. Let us give a hand-wavy example. Acharya et al. (2015) construct a procedure to test $H_{0}:P^{\text{true}}\in\mathcal{M}(\underline{P}_{1})$ versus $H_{1}:P^{\text{true}}\not\in\mathcal{M}(\underline{P}_{1})$. To carry out the test, they first split the available sample size into two parts, and use the first part to “project” $P^{\text{true}}$ onto $\mathcal{M}(\underline{P}_{1})$; denote by $\hat{P}^{\text{true}}$ such a projection. Then, with the second part, they test $H_{0}^{\prime}:P^{\text{true}}\in B_{\eta}(\hat{P}^{\text{true}})$, where $B_{\eta}(\hat{P}^{\text{true}})$ is a ball (with respect to some metric, e.g. the Wasserstein one) of radius $\eta$ centered at $\hat{P}^{\text{true}}$, versus $H_{1}^{\prime}:P^{\text{true}}\not\in B_{\eta}(\hat{P}^{\text{true}})$, and they show that this procedure corresponds to the original test. In the spirit of Ramdas et al. (2017), we conjecture that – if the goal of the researcher is to verify which credal set between $\mathcal{M}(\underline{P}_{1})$ and $\mathcal{M}(\underline{P}_{2})$ contains $P^{\text{true}}$ – this procedure could be extended to one that includes the notion of Wasserstein metric between $\underline{P}_{1}$ and $\underline{P}_{2}$ to take into account how far $\mathcal{M}(\underline{P}_{1})$ and $\mathcal{M}(\underline{P}_{2})$ are from each other, thus making our contributions relevant.
@@
Finally, our findings can be seminal in starting a new field of inquiry that generalizes ergodic transport theory (Lopes and Mengue, 2012; Kabir and Lee, 2020; Chen and Li, 2020) to the credal setting. This may also be related to the field of computer vision (and especially the theory of convolutional autoencoders (Yu et al., 2023)). To see how, we refer the interested reader to Caprio (2024), where the ergodicity of (a version of) imprecise Markov processes is related to the behavior of the outputs of a convolutional autoencoder, as the inputs are perturbed.
@@
Structure of the Paper. The paper is arranged as follows. Section 2 gives the necessary background on credal sets. Our results pertaining the lower probability version of Monge’s and Kantorovich’s problems are presented in Section 3. Section 4 concludes our work.
@@ 2 Background
In this section, we introduce the concepts of the core and of the Choquet integral. The reader who is familiar with these notions can skip to Section 3.
The main tool we work with in this paper is a particular type of a credal set (convex and weak-closed set of probabilities (Levi, 1980)), [^4] The weak convergence is the setwise convergence on the fixed $\sigma$-algebra. that is, what economists and operations researchers call the core (of an exact capacity) (Cerreia-Vioglio et al., 2015; Caprio and Mukherjee, 2023; Miranda and Montes, 2023).
Given a capacity of interest – in this paper, it will always be a lower probability $\underline{P}$, i.e. a set function on the $\sigma$-algebra of interest, mapping in $[0,1]$, which is the lower envelope of a weak-compact set (Cerreia-Vioglio et al., 2015, Section 2.1.(viii)) – on a generic measurable space $(\mathcal{X},\mathcal{F})$, the core is defined as
$\mathcal{M}^{\text{fa}}(\underline{P})\coloneqq\{P\in\Delta^{\text{fa}}_{\mathcal{X}}:P(A)\geq\underline{P}(A)\text{, }\forall A\in\mathcal{F}\},$ (1)
where $\Delta^{\text{fa}}_{\mathcal{X}}$ denotes the set of finitely additive probabilities on $(\mathcal{X},\mathcal{F})$.
We focus on cores for two main reasons. First, in general we have that the convex hull of a finite set of finitely additive probabilities on $\mathcal{X}$ is a proper subset of the core of the lower probability associated with that set, see e.g. (Amarante and Maccheroni, 2006, Example 1) and (Amarante et al., 2006, Examples 6,7,8). That is, given $\{P_{k}\}_{k=1}^{K}\subset\Delta^{\text{fa}}_{\mathcal{X}}$, $K<\infty$, we have that $\text{CH}(\{P_{k}\}_{k=1}^{K})\subset\mathcal{M}^{\text{fa}}(\underline{P})$, where CH denotes the convex hull operator, and $\underline{P}(A)=\inf_{P\in\text{CH}(\{P_{k}\}_{k=1}^{K})}P(A)$, for all $A\in\mathcal{F}$. Hence, focusing on the core gives us more generality.
Second, the core is uniquely identified by its lower probability (Gong and Meng, 2021). To see this, notice that by knowing $\underline{P}$, we can reconstruct the set by simply considering all finitely additive probability measures on $\mathcal{X}$ that set-wise dominate $\underline{P}$.
Before proceeding to the main results of this paper, we need to introduce the concepts of pushforward lower probability (PLP) and of Choquet integral.
@@
Given two measurable spaces $(\mathcal{X},\mathcal{F})$ and $(\mathcal{Y},\mathcal{G})$, a measurable mapping $T:\mathcal{X}\rightarrow\mathcal{Y}$, and a (coherent à la Walley (1991, Section 2.5)) lower probability $\underline{P}=\inf_{P\in\mathcal{M}^{\text{fa}}(\underline{P})}P$, the pushforward of $\underline{P}$ is the (set) function $T_{\#}\underline{P}:\mathcal{G}\rightarrow[0,1]$ such that
@@
Lemma 2 – whose proof relates PLPs to the mathematics of ambiguity (Marinacci and Montrucchio, 2004), and is postponed to Appendix C – entails that $T_{\#}\underline{P}$ is a coherent lower probability on $\mathcal{G}$, hence a superadditive version of a pushforward probability measure (Walley, 1991, Section 1.6.4). We now introduce Choquet integrals (Choquet, 1954), (Troffaes and de Cooman, 2014, Section C.2).
@@
$\begin{split}f\mapsto\int_{\mathcal{Z}}f\text{d}\underline{P}&\coloneqq\int_{0}^{\infty}\underline{P}^{\star}(\{f^{+}\geq t\})\text{d}t\\ &-\int_{0}^{\infty}\left[\underline{P}^{\star}(\mathcal{Z})-\underline{P}^{\star}(\{f^{-}\geq t\}^{c})\right]\text{d}t\end{split}$ (2)
called the Choquet integral of $f$ with respect to $\underline{P}$, provided that the difference on the right-hand side is well defined. There, $f^{+}\coloneqq 0\vee f$, and $f^{-}\coloneqq-(0\wedge f)$. Also, $\underline{P}^{\star}(A)\coloneqq\sup_{B\subseteq A}\underline{P}(B)$, for all $A\in\mathcal{H}$, is the inner lower probability (Walley, 1991, Chapter 3.1), [^5] We need to work with $\underline{P}^{\star}$ because $f$ may not be measurable. When it is, $\underline{P}^{\star}=\underline{P}$. and the integrals are (improper) Riemannian integrals. [^6] For a primer on Riemannian integrals, see Troffaes and de Cooman (2014, Section C.1).
@@
Using the same notation as Definition 2, suppose the Choquet integral $\int_{\mathcal{Z}}f\text{\em\text{d}}\underline{P}$ of $f$ with respect to $\underline{P}$ exists. Then,
$\int_{\mathcal{Z}}f\text{\em\text{d}}\underline{P}$ $=\int_{0}^{\infty}\underline{P}^{\star}(\{f\geq t\})\text{\em\text{d}}t$
$+\int_{-\infty}^{0}\left[\underline{P}^{\star}(\{f\geq t\})-\underline{P}^{\star}(\mathcal{Z})\right]\text{\em\text{d}}t.$
In addition, if $f$ is bounded or Borel measurable, then it is Choquet integrable with respect to $\underline{P}$, that is, its Choquet integral $\int_{\mathcal{Z}}f\text{\em\text{d}}\underline{P}$ exists.
Corollary 4.1 (A Simplification of the Choquet Integral).
Using the same notation as Definition 2, if $f$ is positive and measurable, then $\int_{\mathcal{Z}}f\text{\em\text{d}}\underline{P}=\int_{0}^{\infty}\underline{P}(\{f\geq t\})\text{\em\text{d}}t$. If $f$ is also bounded, then the weak inequality can be substituted by a strict one.
Proof.
The first part of the statement comes from Proposition 4 and Marinacci and Montrucchio (2004, Equation (11)). The second part is a consequence of Marinacci and Montrucchio (2004, Proposition 17). A similar result to Corollary 4.1 was proven by Grabisch (2016). ∎